OpenStudy (shrek_wazowski):

-==||WILL FAN AND MEDAL||==- Solve the equation. x - 2/7 = 3 A. x = -3 2/7 B. x = -2 5/7 C. x = 2 5/7 D. x = 3 5/7

OpenStudy (mathmale):

Pls ensure you've copied the answer choices correctly. One of them appears to be the correct answer to this problem, except for its sign.

OpenStudy (shrek_wazowski):

I've copied them correctly.. Which one has the wrong sign?

OpenStudy (shrek_wazowski):

@mathmale

OpenStudy (ibrafacts):

thats the one that should be right abut u mess up the sign

OpenStudy (ibrafacts):

3 should not be neg u didnt move it at all

OpenStudy (shrek_wazowski):

The sign says negative on my question..

OpenStudy (ibrafacts):

then idk

OpenStudy (mathmale):

Solve the equation. x - 2/7 = 3. Step 1: Add the fraction 2/7 to both sides, and simplify your results.

OpenStudy (shrek_wazowski):

Solve the equation. x – 2/7 = 3 A. x = –3 2/7 B. x = –2 5/7 C. x = 2 5/7 D. x = 3 2/7

OpenStudy (ibrafacts):

yes u see u did copy it wrong

OpenStudy (ibrafacts):

i see the right answer already

OpenStudy (shrek_wazowski):

so it should be A. x = +3 2/7 ????

OpenStudy (ibrafacts):

yes

OpenStudy (mathmale):

Yes.

OpenStudy (shrek_wazowski):

So could it be D. ??

OpenStudy (ibrafacts):

its not could it is

OpenStudy (shrek_wazowski):

Okay :) Thanks alot!

OpenStudy (ibrafacts):

yup:)